The average daily production of a certain product in a plant is recorded. Registration was carried out for 54 days and an average of 1,350 liters with a standard deviation of 260 liters was obtained. Test the hypothesis at a significance level of 5%, that the average daily production is 1400 liters.

A) z = -1,413> -1.96 Ho is not rejected

B) z = -1,413> -1,645 Ho is not rejected

C) z = -1,413> -1,645 Ho is rejected

D) z = -1.413> -1.96 Ho is rejected
Answer #1

Test statistic = z

= (\bar x - \mu ) / \sigma    / \sqrt n

= (1350 - 1400) / 260 / \sqrt 54

= -1.413

Test statistic = -1.413

A) z = -1,413> -1.96 Ho is not rejected